OpenStudy (anonymous):

If f(x) = x − 4 and g(x) = 1 over 2x − 2, find f[g(−6)]. please help (IN A HURRY)

OpenStudy (anonymous):

@ganeshie8 @mathstudent55 @jdoe0001

jigglypuff314 (jigglypuff314):

first plug in -6 for the x values in the function of g(x)

jigglypuff314 (jigglypuff314):

then simplify that then plug in what you got into the x values in the function of f(x)

OpenStudy (anonymous):

:o

jigglypuff314 (jigglypuff314):

get it?

OpenStudy (anonymous):

how to i plug in the -6 in the fraction

OpenStudy (anonymous):

do*

jigglypuff314 (jigglypuff314):

\[\frac{ 1 }{ 2x-2 } \rightarrow \frac{ 1 }{ 2(-6)-2 } \rightarrow \frac{ 1 }{ -12-2 } \rightarrow \frac{ 1 }{ -14 }\]

OpenStudy (anonymous):

ok so i got -10 on the left side and 1 over -14 on the right side what do i do next

jigglypuff314 (jigglypuff314):

no, f(g(-6)) means to plug in -6 into the function of g(x) and take what you got from that, and plug that into the function of f(x)

jigglypuff314 (jigglypuff314):

so you plug in 1 over -14 into f(x)

OpenStudy (anonymous):

this is confusing @jigglypuff314

jigglypuff314 (jigglypuff314):

g(-6) means to plug in -6 into the function of g(x) f (g(-6)) means to plug in what g(-6) equals into the function of f(x)

OpenStudy (anonymous):

>.< can you please like write out the problem for me step by step

jigglypuff314 (jigglypuff314):

\[g(x) = \frac{ 1 }{ 2x-2 } \rightarrow g(-6)=\frac{ 1 }{ 2(-6)-2 }\] do you understand this so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you plugged in the -6

jigglypuff314 (jigglypuff314):

then you simplify that

OpenStudy (anonymous):

do i multiply 2 and negative 6

jigglypuff314 (jigglypuff314):

yes

OpenStudy (anonymous):

and then - 12-2

OpenStudy (anonymous):

is it 1 over negative 14

jigglypuff314 (jigglypuff314):

yes :)

OpenStudy (anonymous):

thats it?

jigglypuff314 (jigglypuff314):

then you plug 1 over 14 into x-4

OpenStudy (anonymous):

how do i do that

jigglypuff314 (jigglypuff314):

so 1/14 - 4

OpenStudy (anonymous):

14-4

jigglypuff314 (jigglypuff314):

no, (1 over -14) - 4

OpenStudy (anonymous):

i multiply - 4

jigglypuff314 (jigglypuff314):

no, 4 = 56 over 14 so 1 over negative 14 minus 56 over 14

OpenStudy (anonymous):

ho do i solve that

OpenStudy (anonymous):

55over -28

jigglypuff314 (jigglypuff314):

no, negative number minus number equals negative added numbers and when adding fractions (or numbers over numbers) you don't add the bottom two numbers.

OpenStudy (anonymous):

so 55 over 14?

jigglypuff314 (jigglypuff314):

-1-56 = -57 so your answer is negative 57 over 14

OpenStudy (anonymous):

thats not right

jigglypuff314 (jigglypuff314):

how is it not?

OpenStudy (anonymous):

my choices are 9 3 −9 −3

OpenStudy (anonymous):

soo ?

jigglypuff314 (jigglypuff314):

hmm. i don't know...

OpenStudy (anonymous):

:o

OpenStudy (anonymous):

come on! your a pokemon!

OpenStudy (anonymous):

help!! we been on this question for 33 minutes i really need someone to help me

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

@Easyaspi314 @shamil98